# The NordDRG AI Benchmark for Large Language Models

### v1

@@ 1.2 Contributions
We address this shortfall by releasing NordDRG-AI-Benchmark, the first public benchmark that unites clinical coding, multilingual reasoning and health-care finance. Specifically, we contribute (i) a machine-readable dump of the Nordic DRG definition tables, (ii) governance manuals that capture real-world change-management workflows, and (iii) 14 task-oriented prompts spanning code lookup, cross-table inference and qualitative audit scenarios. All artefacts are released to catalyse research on transparent, regulator-compliant language models for hospital funding.
@@ 1.5 Research objectives
    O2 — Task operationalisation.  Instantiate a balanced test-suite of 14 CaseMix prompts (9 quantitative, 5 qualitative) that reflect real workflows of clinicians, coders, and policy analysts, together with a gold answer key for automatic or rubric-based scoring.
@@ 2.1 NordDRG CaseMix System
NordDRG is distributed as a set of definition tables—about twenty inter-linked sheets released annually in Excel, CSV, and JSON formats—that encode the logic connecting diagnoses, procedures, age/sex splits, cost weights, and national activation flags to individual DRG codes. These tables constitute the authoritative rule base for coding software, technical-change management, and policy analysis.
@@ 2.3 LLM Benchmarks and Leaderboards
Benchmark suites and public leaderboards now play a decisive role in charting progress across the rapidly evolving LLM landscape. Early task collections such as GLUE [Wang et al., 2018] and SuperGLUE [Wang et al., 2019] have given way to multi-domain, multi-skill batteries—e.g. BIG-bench  [bench Collaboration, 2022], MMLU [Hendrycks et al., 2021], and the BabyLM Challenge [Warstadt et al., 2023]—that probe factual knowledge, mathematical reasoning, multilingual transfer, and instruction following. Complementary meta-evaluations such as HELM [Liang et al., 2022] standardise experimental settings and report robustness, bias, and efficiency metrics side by side.
Results from these test suites are aggregated on widely watched leaderboards. The HuggingFace Open LLM Leaderboard publishes automatic scores (e.g. MMLU, ARC, HellaSwag) for hundreds of open-weight models, while the LMSys Chatbot Arena [^2] https://chat.lmsys.org/ crowdsources pairwise human preferences to rank dialogue quality. Other platforms—including Vellum-AI [^3] https://www.vellum.ai/llm-leaderboard and MT-Bench [Zheng et al., 2023]—focus on application-specific criteria such as tool use or code generation. Figure 1 shows a representative snapshot of one such leaderboard, underscoring how transparent, regularly updated scorecards have become the de-facto currency for comparing models, spurring innovation, and diagnosing remaining failure modes.
@@ 3 Design & Development: NordDRG Benchmark
    Procedural comprehension of the change-management workflows that drive yearly tariff updates across six Nordic health systems.
@@
All resources are released under at https://github.com/longshoreforrest/norddrg-ai-benchmark and are grouped into four artefact classes:
  • A1 — Full definition corpus (.xlsx/.csv): roughly twenty relational sheets that encode the complete NordDRG grouper logic—diagnoses, NCSP procedures, age/sex splits, country-activation flags, and surgery properties.
@@
  • identifying DRGs for a specific procedure (e.g. appendectomy),
@@
Nine tasks (IDs 1–9) are quantitative: they admit deterministic answer keys and therefore support exact-match scoring. The remaining five tasks (IDs 10–14) are qualitative: they require free-text justifications or change-request drafts and are graded with a rubric-based human review.
Taken together, the fourteen prompts constitute a self-contained test suite for head-to-head comparisons of prompt engineering, retrieval augmentation, and fine-tuning strategies. By releasing this benchmark, we provide the first public, reproducible yardstick for LLMs in hospital-funding contexts, thereby closing the application and evaluation gaps identified earlier and laying the groundwork for transparent, regulator-compliant clinical language models.
@@ 3.1 Artefact: Definition Tables
  • drg_logic — rules that map diagnosis, procedure, age/sex limits, CC/MCC status, and surgery flags to a unique DRG code; ordered by the ORD execution column.
  • dg_feat / proc_feat — full ICD-10 and NCSP catalogues annotated with properties referenced by drg_logic (e.g. DGCAT, COMPL, OR, PROCPRO).
@@
  • age_split and country_flag tables that hold paediatric/geriatric break-points and activation flags for each Nordic version.
Each worksheet exposes stable identifier columns (e.g. DRG, ICD, NCSP, ORD), allowing the tables to be joined loss-lessly and extended by appending new rows in future releases. A representative excerpt is shown in Figure 2.
Figure 2: Materials used: NordDRG definition tables https://nordcase.org/
@@ 3.2 Artefact: Expert documentation
Figure 3: Materials used: NordDRG definition table expert documentation https://nordcase.org/
@@ 3.3 Artefact: Definition Tables: Part of the tables FI
Artefact 3.3 distils the full workbook down to a single, lightweight file. The sheet contains only those DRG codes that are activated in the Finnish version and provides their official Finnish descriptions (DRG_TEXT_FI). Auxiliary columns preserve the DRG code, MDC, weight class and version stamp so that rows can still be cross-referenced against the master tables.
@@ 3.7 Artefact: NordDRG CaseMix questions
Artefact 3.7 contains the 14 CaseMix questions that form the core of the benchmark. The benchmark ships with a compact task suite that mirrors day-to-day queries posed by clinicians, health-economists and DRG coders. Table 3 lists the nine quantitative prompts (IDs 1–9). Their answers are deterministic code sets, so an LLM can be scored automatically by exact match against the ground-truth sheet. Table 4 adds five qualitative prompts (IDs 10–14) that demand free-text justifications—e.g. drafting a technical-change request graded with a human review.
@@ 8 Diagnosis Properties Medium Which diagnosis (ICD) codes are connected to diagnosis property 14X03?
12 Quality Assurance Hard Read the names for the ICD codes already linked to 14X03. Then review all diagnosis codes in the dg sheet. Which five codes are most likely missing from 14X03 based on their text descriptions?

### v2

@@
All artefacts are released at GitHub  [^1] https://github.com/longshoreforrest/norddrg-ai-benchmark.
@@
Keywords Artificial Intelligence AI, Large Language Models LLM, Diagnosis Related Groups DRG, NordDRG, CaseMix systems, Clinical classifications
@@ 1.2 Contributions
We address this shortfall by releasing NordDRG-AI-Benchmark, the first public benchmark that unites clinical coding, multilingual reasoning and health-care finance. Specifically, we contribute (i) a machine-readable dump of the NordDRG definition tables, (ii) governance manuals that capture real-world change-management workflows, and (iii) 14 task-oriented prompts spanning code lookup, cross-table inference and qualitative audit scenarios. All artefacts are released to catalyse research on transparent, regulator-compliant language models for hospital funding.
@@ 1.5 Research objectives
    O2 — Task operationalisation.  Instantiate a balanced test-suite of 14 DRG CaseMix prompts (9 quantitative, 5 qualitative) that reflect real workflows of clinicians, coders, and policy analysts, together with a gold answer key for automatic or rubric-based scoring.
@@ 2.1 NordDRG CaseMix System
NordDRG is distributed as a set of definition tables—about twenty inter-linked sheets released annually in Excel, CSV, and JSON formats—that encode the logic connecting diagnosis and procedure codes, age/sex splits, and national activation flags to individual DRG codes. These tables constitute the authoritative rule base for coding software, technical-change management, and policy analysis.
@@ 2.3 LLM Benchmarks and Leaderboards
Benchmark suites and public leaderboards now play a decisive role in charting progress across the rapidly evolving LLM landscape. Early task collections such as GLUE [Wang et al., 2018] and SuperGLUE [Wang et al., 2019] have given way to multi-domain, multi-skill batteries—e.g., BIG-bench, [bench Collaboration, 2022], MMLU [Hendrycks et al., 2021], and the BabyLM Challenge [Warstadt et al., 2023]—that probe factual knowledge, mathematical reasoning, multilingual transfer, and instruction following. Complementary meta-evaluations such as HELM [Liang et al., 2022] standardise experimental settings and report robustness, bias, and efficiency metrics side by side.
Results from these test suites are aggregated on widely watched leaderboards. The HuggingFace Open LLM Leaderboard publishes automatic scores (e.g., MMLU, ARC, HellaSwag) for hundreds of open-weight models, while the LMSys Chatbot Arena [^3] https://chat.lmsys.org/ crowdsources pairwise human preferences to rank dialogue quality. Other platforms—including Vellum-AI [^4] https://www.vellum.ai/llm-leaderboard and MT-Bench [Zheng et al., 2023]—focus on application-specific criteria such as tool use or code generation. Figure 1 shows a representative snapshot of one such leaderboard, underscoring how transparent, regularly updated scorecards have become the de-facto currency for comparing models, spurring innovation, and diagnosing remaining failure modes.
@@ 3 Design & Development: NordDRG Benchmark
    Procedural comprehension of the change-management workflows that drive yearly tariff updates across six NordDRG user countries.
@@
All resources are released under at GitHub  [^5] https://github.com/longshoreforrest/norddrg-ai-benchmark and are grouped into four artefact classes:
  • A1 — Full definition corpus (.xlsx/.csv): roughly twenty relational sheets that encode the complete NordDRG grouper logic—diagnoses, NCSP procedures, age/sex splits, country-activation flags, surgery properties and grouping features.
@@
  • identifying DRGs for a specific procedure (e.g., appendectomy),
@@
Nine grouping feature tasks (IDs 1–9) are quantitative: they admit deterministic answer keys and therefore support exact-match scoring. The remaining five tasks (IDs 10–14) are qualitative: they require free-text justifications or change-request drafts and are graded with a rubric-based human review.
Taken together, the fourteen prompts constitute a self-contained test suite for head-to-head comparisons of prompt engineering, retrieval augmentation, and fine-tuning strategies. By releasing this benchmark, we provide the first public, reproducible yardstick for LLMs in hospital-funding contexts, thereby closing partially the application and evaluation gaps identified earlier and laying the groundwork for transparent, regulator-compliant clinical language models.
@@ 3.1 Artefact: Definition Tables
  • drg_logic — rules that map diagnosis, procedure, age limits, CC/MCC status, and surgery flags to a unique DRG code; ordered by the ORD execution column.
  • dg_feat / proc_feat — full ICD-10 and NCSP diagnosis and procedure code catalogues annotated with properties referenced by drg_logic (e.g. DGCAT, COMPL, OR, PROCPRO).
@@
  • age_split and country_flag tables that hold paediatric/geriatric break-points and activation flags for each national version.
Each worksheet exposes stable identifier columns (e.g., DRG, ICD, NCSP, ORD), allowing the tables to be joined loss-lessly and extended by appending new rows in future releases. A representative excerpt is shown in Figure 2.
Figure 2: Materials used: NordDRG definition tables https://nordcase.org/forum/
@@ 3.2 Artefact: Expert documentation
Figure 3: Materials used: NordDRG definition table expert documentation https://nordcase.org/forum/
@@ 3.3 Artefact: Definition Tables: Part of the tables FI
Artefact 3.3 distils the full workbook down to a single, lightweight file. The sheet contains only those DRG codes that are activated in the Finnish version and provides their official Finnish descriptions (DRG_TEXT_FI). Auxiliary columns preserve the DRG code, MDC and version stamp so that rows can still be cross-referenced against the master tables.
@@ 3.7 Artefact: NordDRG CaseMix questions
Artefact 3.7 contains the 14 CaseMix questions that form the core of the benchmark. The benchmark ships with a compact task suite that mirrors day-to-day queries posed by clinicians, health-economists and DRG coders. Table 3 lists the nine quantitative prompts (IDs 1–9). Their answers are deterministic code sets, so an LLM can be scored automatically by exact match against the ground-truth sheet. Table 4 adds five qualitative prompts (IDs 10–14) that demand free-text justifications—e.g., drafting a technical-change request graded with a human review.
@@ 8 Diagnosis Properties Medium Which diagnosis (ICD) codes are connected to diagnosis property 14X03?
12 Quality Assurance Hard Read the names for the ICD codes already linked to 14X03. Then, review all diagnosis codes in the dg sheet. Which five codes are most likely missing from 14X03 based on their text descriptions?
@@ 6 Communication
Acknowledgements
We would like to acknowledge the Nordic CaseMix Centre (NCC) for its significant contributions to the development of NordDRG-AI-Benchmark by providing access to NordDRG definition tables.
We also thank Kristiina Kahur, CEO of NCC, for her valuable comments on the manuscript, which helped improve the clarity and accuracy of this paper.